# ADAPT: A Pseudo-labeling Approach to Combat Concept Drift in Malware Detection

### v1

@@ I Introduction
The ever-evolving nature of malware poses ongoing challenges to cybersecurity. Machine learning has been widely adopted for malware detection across various platforms, offering significant improvements over traditional signature-based and rule-based methods [13, 8, 49, 52, 36]. However, the effectiveness of these models is predicated on the assumption that the statistical properties of the training data remain stable over time [18]. In practice, this assumption rarely holds due to the dynamic nature of malware and the continuous evolution of underlying platforms, which introduce new features, APIs, and programming practices [38]. This leads to the problem of concept drift—a shift in the underlying data distribution that results in a significant degradation in model performance, necessitating constant monitoring and adaptation [43, 50].
Dealing with concept drift in malware detection is a multifaceted challenge. One approach aims to design inherently more drift-resistant systems by creating more robust feature spaces [49, 77]. However, concept drift remains inevitable due to the dynamic nature of the problem landscape, and it is unclear whether a malware representation immune to concept drift can be developed [15]. Another approach to addressing concept drift is periodic model retraining [46, 50]. Effective retraining first requires detecting when the model has become outdated, which triggers the need for an update [37, 15, 73]. Since retraining requires access to manually labeled drifted samples, active learning has been widely used to reduce the annotation effort [47, 23]. However, obtaining high-quality annotations for malware remains challenging, constituting a significant portion of the overall costs of developing an ML model, especially given the sheer volume of emerging malware variants [45, 19].
@@
Self-training provides an alternative approach for concept drift adaptation without requiring additional annotations beyond the initial training data [7]. Pseudo-labeling is a learning paradigm where the model generates labels for unlabeled data, treating its predictions as ground truth to train itself [40, 57] further. A notable application of this approach for malware detection is DroidEvolver [70], which employs pseudo-labeling for Android malware detection by using an ensemble of online learning algorithms to eliminate labeling costs after the initial training phase. However, a later study [38] revealed that when experimental biases, such as unrealistic benign-to-malware ratios in the dataset, are addressed, DroidEvolver’s performance deteriorates significantly due to confirmation bias [12] and a catastrophic self-poisoning effect, leading to a sudden and severe performance drop. While the work in [38] proposed an enhanced version of DroidEvolver to mitigate these issues, the performance still deteriorates rapidly after a short period when solely relying on pseudo-labels. Thus, designing an effective pseudo-labeling-based approach for malware detection under concept drift remains a significant challenge.
@@
In this work, we propose ADAPT: Adaptive Drift-aware Algorithm using Pseudo-labeling for malware detecTion, a semi-supervised learning approach tailored for concept drift in malware detection. Unlike general-purpose methods, ADAPT accounts for the asymmetric nature of drift, which predominantly affects malware rather than benign samples [15, 26]. We introduce a drift-aware pseudo-labeling strategy that selectively propagates labels based on class-specific drift behavior. Data augmentation plays a crucial role in the success of semi-supervised learning algorithms [59, 17]. While well-established in domains such as image classification [27], data augmentation techniques are relatively rare in malware classification. Nevertheless, prior work has shown promising results using data augmentation in noisy label settings for malware classification [69]. We extend this approach to the concept drift scenario, where the model is periodically retrained. To further improve robustness, we incorporate mixup regularization [74, 64], which helps mitigate self-poisoning by improving confidence calibration. This is particularly important for our pseudo-labeling strategy, as it ensures that predicted probabilities reflect the actual likelihood of correctness more accurately.
@@
    We introduce ADAPT, a semi-supervised learning algorithm tailored for concept drift in malware detection. It mitigates catastrophic self-poisoning through drift-aware pseudo-labeling.
@@ II-B Concept Drift
Concept drift in malware detection arises from both benign and malicious changes. While API updates and new functionalities in benign apps contribute to drift, the primary driver is the evolving behavior of malware [25, 77]. Adversaries continually adapt through obfuscation, packing, and novel attack strategies to evade detection [1, 9, 62]. These shifts alter the distribution of malicious behavior between training and testing data [15], leading to degraded classifier performance and reduced detection rates over time [5, 50, 37].
@@ II-C Concept Drift Adaptation
Approaches like Transcend [37] and Transcendent [15] tackle drift through classification with rejection, using conformal evaluators to flag uncertain samples for expert review. However, the reliance on manual labeling limits retraining frequency due to cost and resource constraints [45, 70]. To address this, active learning has been used to reduce annotation effort by selectively labeling uncertain samples [72, 23], with studies showing up to 8x reduction in labeling cost [23]. More recently, pseudo-labeling has emerged as a promising alternative, enabling model updates without new annotations [70, 38]. Our research focuses on this setting; however, we also consider the active learning scenario in Section VIII.
@@ III Research Motivation
Figure 1: Impact of covariate shift $p(x)$ and concept shift $p(y|x)$ on the first year of the test set for the two Android datasets, evaluated using Optimal Transport Dataset Distance and Fréchet Distance.
@@
While it is not feasible to precisely measure different forms of dataset shift from a finite set of samples, these shifts can be approximated using various metrics. In this work, we use the Optimal Transport Dataset Distance (OTDD) [6] to approximate covariate shift and the Fréchet Dataset Distance (FDD) [35] to approximate concept shift, following the methodology suggested in [31].
OTDD [6] measures the distance between two distributions by computing the minimum cost of transforming one distribution into another, leveraging optimal transport theory. Using a Gaussian approximation, we calculate the OTDD between the first year of testing data and the original training data. This distance is computed directly in the feature space without relying on a trained model.
Fréchet Dataset Distance (FDD) [35], initially used in generative models, measures the similarity between two distributions by comparing their intermediate feature representations in a reference classifier. To compute the FDD, we use the MLP model trained with the best hyperparameters (as described in Section VI-C). Specifically, we take the penultimate layer features of the trained model for each month’s samples and compare them with the features from the training month to obtain the FDD. More details on these metrics can be found in their respective papers [6, 35].
The results for the two Android malware detection datasets used in this study are shown in Figure 1. The magnitude of both covariate and concept shift is higher in the Drebin [13] dataset compared to APIGraph [77], as APIGraph employs a more compact feature space that captures semantically similar API calls. The general trend shows that these shifts gradually increase in both datasets. In particular, these shifts lead to poorer malware detection performance when the model is not updated. This highlights that malware detection datasets often exhibit different shifts, which can interact in complex ways to degrade classifier performance.
@@
Figure 2: Adaptation to distribution shift for a binary classifier. The solid line represents the original classifier, and the dotted lines represent the adapted classifier trained with pseudo-labels.
Self-training for drift adaptation
@@
Self-training with pseudo-labeling can allow a model to adapt to evolving data distributions without requiring continuous manual labeling [39]. This process is illustrated in Figure 2 using a 2D toy dataset with a logistic regression classifier. In this example, both $p(x)$ and $p(y|x)$ gradually change from $t=0$ to $t=3$, with the dataset undergoing continuous rotation. As a result, the classifier’s initial decision boundary (solid line) misclassifies most samples by $t=3$.
@@
Nevertheless, malware datasets typically undergo gradual distributional shifts, as illustrated in Figure 1, where both covariate and concept shifts increase over time. Based on this observation, we hypothesize that an effective self-training approach can continuously adapt the model to account for these evolving shifts. Notably, prior work by Kumar et al. [39] establishes formal error bounds of self-training under the assumption of gradual distributional shifts. Our method builds on this foundation by extending their framework, and we provide a rigorous theoretical analysis in Appendix K.
@@ IV-A Adaptive Drift-Aware Pseudo-Labeling
Figure 3: Mean and standard deviation of predicted probabilities for benign and malware classes using XGBoost and MLP models on two datasets. Malware predictions typically have lower confidence, indicating concept drift primarily affects the malware class.
@@ IV-B Data Augmentation with Label Consistency
$\{[0,1,0],[1,1,0],[0,0,1],[0,1,1]\}.$
The value of the second feature of $\hat{\mathbf{x}}$ would be randomly sampled from the set $\{1,1,0,1\}$, which are the observed values of the second feature across all samples.
Figure 4: Percentage of samples in the Drebin and APIGraph datasets where augmentation causes the model’s originally correct prediction to flip (i.e., benign to malicious or malicious to benign).
@@
To address this, we propose two key modifications to ensure label consistency during augmentation. First, when replacing a masked feature, we restrict the replacement values only from samples of the same class. Although this setting was considered in [69], we argue that preserving label consistency is even more critical in our iterative training framework. Second, after the initial training phase, we leverage the current model to predict the label of the augmented sample. The augmented sample is included in the dataset only if the model’s prediction matches the original label to improve label consistency. However, it is important to note that feature consistency is not explicitly enforced during augmentation. This means the generated features may not always represent valid samples, potentially leading to inconsistencies. We explore this issue in more detail in Appendix J.
@@ IV-C Confidence Calibration with Mixup
However, this approach is not applicable to models like Random Forests, which do not support fractional targets for classification tasks. To handle this, when mixing two samples with different labels, we assign the label corresponding to the sample with the higher interpolation coefficient.
Figure 5: APIGraph dataset
We demonstrate the effectiveness of Mixup in improving model calibration in Figure 6 using the XGBoost model on the APIGraph dataset. The scatter plots illustrate the model’s accuracy versus its confidence for different confidence bins across three models: the base model trained without data augmentation or Mixup (left), the model trained with only data augmentation (middle), and the model trained with both Mixup and data augmentation (right), all evaluated on the validation dataset. A well-calibrated model will have most points close to the $y=x$ line, indicating that the model’s confidence aligns with its actual accuracy. The figure shows that the base model exhibits significant overconfidence, where its confidence exceeds its true accuracy. This overconfidence can be particularly problematic when using thresholds to select pseudo-labeled samples. In contrast, the model trained with data augmentation shows improved calibration, and the model trained with both Mixup and data augmentation further enhances this, with its confidence more accurately reflecting its accuracy, especially in high-confidence regions.
@@ VI-B Self-Training Baselines
We compare our method against four self-training baselines: Adaptive Random Forest (ARF) [32], DroidEvolver++ (DE++) [38], Insomnia [10], and MORSE [69]. ARF and DE++ are online learning methods that address concept drift without requiring access to the original training dataset. In contrast, Insomnia and MORSE follow a semi-supervised paradigm, leveraging the original training data and pseudo-labeled samples for model updates. While our approach (ADAPT) also assumes access to the original training data, we additionally explore scenarios where this access is restricted. Further implementation details for all baseline methods are provided in Appendix B, and the restricted setting without original training data is discussed in Appendix G.
@@ VI-C Hyperparameter Tuning
As shown in previous works [23, 2], continuous learning settings require their own hyperparameter tuning. In our experiments, we perform joint hyperparameter tuning for the machine learning models and our proposed method using the validation set.
We introduce five hyperparameters in ADAPT: the benign base threshold $\tau_{b}$, the malware base threshold $\tau_{m}$, the threshold weight $\lambda$ for adaptive thresholding, the masking probability for augmentation $p_{a}$, and the Mixup parameter $\alpha$. Each baseline self training method has its own set of hyperparameters. Specifically:
  • ARF: threshold $\tau$
  • DE++: app buffer size $B$, the ratio of malware to benign samples in the buffer $m_{r}$, and the threshold for model aging $\tau_{a}$
  • Insomnia: the fraction of samples considered for pseudo-labeling $p_{r}$
  • MORSE: threshold $\tau$, weak augmentation masking probability $p_{w}$, and strong augmentation masking probability $p_{s}$
A detailed list of model and self-training hyperparameters, along with their respective ranges, is provided in the Appendix L. Since different methods introduce varying numbers of hyperparameters, we use a fixed hyperparameter search budget to ensure a fair comparison between methods, as recommended in [2]. Specifically, we conduct 200 random hyperparameter searches [16] for each method on each dataset.

### v2

@@ I Introduction
The ever-evolving nature of malware poses ongoing challenges to cybersecurity. While machine learning offers substantial improvements over traditional signature- and rule-based detection [12, 7, 54, 57, 39], its effectiveness depends on the assumption that data distributions remain stable [18]. In reality, continuous changes in malware and underlying platforms—such as new features, APIs, and programming practices [41]—frequently violate this assumption, causing concept drift and leading to significant model degradation unless adaptation measures are taken [47, 55].
Dealing with concept drift in malware detection is a multifaceted challenge. While some approaches focus on building more robust feature spaces to resist drift [54, 83], true immunity remains elusive given malware’s dynamic landscape [14]. Periodic model retraining offers another solution [50, 55], but requires timely drift detection and access to newly labeled samples [40, 14, 80]. Active learning helps reduce annotation effort [51, 24], yet obtaining high-quality malware labels is costly due to the sheer volume of emerging malware variants [49, 20].
@@
Self-training offers a way to adapt to concept drift without requiring new annotations beyond the initial training data [6]. Pseudo-labeling is a learning paradigm where the model generates labels for unlabeled data, treating its predictions as ground truth to train itself [44, 62] further. A notable application of this approach for malware detection is DroidEvolver [77], which employs pseudo-labeling for Android malware detection by using an ensemble of online learning algorithms to eliminate labeling costs after the initial training phase. However, a later study [41] revealed that when experimental biases, such as unrealistic benign-to-malware ratios in the dataset, are addressed, DroidEvolver’s performance deteriorates significantly due to confirmation bias [11] and a catastrophic self-poisoning effect, leading to a sudden and severe performance drop. Although the work in [41] proposed an enhanced version of DroidEvolver to mitigate these issues, performance still deteriorates rapidly after a short period when solely relying on pseudo-labels. Thus, designing a practical pseudo-labeling-based approach for malware detection under concept drift remains a significant challenge.
@@
In this work, we propose ADAPT: Adaptive Drift-aware Algorithm using Pseudo-labeling for malware detecTion, a semi-supervised approach tailored for concept drift in malware detection. Unlike general-purpose methods, ADAPT accounts for the asymmetric nature of drift, which more frequently impacts malware than benign samples [14, 27], by introducing a drift-aware pseudo-labeling strategy that selectively propagates labels based on class-specific drift behavior. Data augmentation, though standard in image classification [28], remains uncommon in malware detection, but has shown promise in noisy label settings [76]. We extend these augmentation strategies to the concept drift scenario, periodically retraining the model to adapt to changing distributions. To further enhance robustness, we incorporate mixup regularization [81, 70], which improves confidence calibration and helps mitigate the self-poisoning effect commonly seen in pseudo-labeling. This ensures that the model’s predicted probabilities better reflect the actual likelihood of correctness, making pseudo-labeling more effective in practice.
@@
    We introduce ADAPT [^1] https://github.com/aiforsec/ADAPT, a semi-supervised learning algorithm tailored for concept drift in malware detection. It mitigates catastrophic self-poisoning through drift-aware pseudo-labeling.
@@ II-B Concept Drift
Concept drift in malware detection arises from both benign and malicious changes. Updates to APIs and the introduction of new functionalities in benign applications contribute to changes in the input feature space, often resulting in what is referred to as virtual drift [63]. In contrast, the primary driver of concept drift is the continually evolving behavior of malware [26, 83], which leads to real concept drift, a shift in the relationship between features and labels [63]. Adversaries regularly adapt through obfuscation, packing, and novel attack strategies to evade detection [1, 8, 68]. These dynamics alter the distribution of malicious behavior between training and testing data [14], degrading classifier performance and reducing detection rates over time [55, 40].
@@ II-C Concept Drift Adaptation
Approaches such as Transcend [40] and Transcendent [14] address concept drift by employing classification with rejection, using conformal evaluators to flag uncertain samples for expert review. However, reliance on manual labeling inherently limits the frequency of model retraining due to cost and resource constraints [49, 77]. Active learning has been proposed to alleviate the annotation burden by selectively labeling only the most uncertain samples [79, 24], with recent studies demonstrating up to an eightfold reduction in labeling costs [24].
Despite these advances, practical challenges such as label delays can significantly increase user exposure to undetected malware [19]. As a result, there is a clear need for interim solutions that can adapt to drift in the absence of timely ground-truth labels. Pseudo-labeling provides one such solution, enabling models to update themselves using confident predictions in place of manual annotations [19, 77, 41].
@@ III Research Motivation
Figure 1: Covariate shift $p(x)$ and concept shift $p(y|x)$ measured by Optimal Transport Dataset Distance (OTDD) and Fréchet Dataset Distance (FDD) across the first year of testing on two Android datasets (Drebin and APIGraph). The corresponding monthly F1-scores from three models (Random Forest, XGBoost, and MLP) are shown to illustrate the relationship between distributional shifts and malware detection performance.
@@
Machine learning models used for malware detection experience degraded performance over time due to concept drift. Although precise measurement of dataset shifts from finite samples is challenging, they can be approximated using metrics such as the Optimal Transport Dataset Distance (OTDD) [5] for covariate shift and the Fréchet Dataset Distance (FDD) [38] for concept shift [33].
OTDD [5] quantifies covariate shift by calculating the minimal cost to transform one data distribution into another using optimal transport theory. Specifically, we approximate this distance with a Gaussian approximation computed directly in the feature space between the training data and each subsequent month’s testing data. In contrast, Fréchet Dataset Distance (FDD) [38], initially introduced in the context of generative modeling, captures concept shift by evaluating the similarity of intermediate feature representations extracted from a trained reference classifier. Here, we employ a Multilayer Perceptron (MLP) model trained with optimal hyperparameters (as detailed in Section VI-C) and compare penultimate layer features from each test month’s samples against those of the training set.
Applying these metrics to two prominent Android malware datasets, Drebin [12] and APIGraph [83] (Figure 1), we observe substantial and progressively increasing covariate and concept shifts. These distributional shifts are notably more pronounced in Drebin, which utilizes a larger, less semantically compact feature space compared to APIGraph’s more concise API-based representation. Although both datasets experience degradation in malware detection accuracy over time, this deterioration more consistently correlates with increased OTDD and FDD distances in APIGraph, highlighting complex interactions between dataset shifts and classifier performance. These insights underscore the critical importance of continuously adapting models to mitigate the negative impacts of concept drift.
@@
Figure 2: Adaptation to distribution shift using pseudo-labels. Solid line: original classifier; dotted lines: adapted classifier trained with pseudo-labels.
Challenges of Label Delays in Practice
Effective adaptation to concept drift fundamentally relies on the availability of timely ground-truth labels for evaluating model predictions. However, in practice, these labels often arrive after significant delays due to resource-intensive processes. Sandbox analyses, essential for automated malware evaluation, face scalability constraints relative to the volume of daily submissions, resulting in prolonged analysis queues [30]. Human analysts are even scarcer, causing manual labeling to become a severe bottleneck with substantial waiting periods [15]. Consequently, these delays open extended vulnerability windows, significantly increasing user exposure to emerging threats and degrading detection performance [19].
Pseudo-labeling for Effective Drift Adaptation
@@
To mitigate these practical challenges associated with label delays, pseudo-labeling emerges as a promising strategy. Pseudo-labeling is a self-training technique in which model-generated predictions serve as provisional labels for unlabeled data, enabling continuous model adaptation without immediate reliance on manual annotations [43]. This process is illustrated in Figure 2 using a 2D toy dataset with a logistic regression classifier. In this example, both $p(x)$ and $p(y|x)$ gradually change from $t=0$ to $t=3$, with the dataset undergoing continuous rotation. As a result, the classifier’s initial decision boundary (solid line) misclassifies most samples by $t=3$.
@@
Nevertheless, malware datasets typically undergo gradual distributional shifts, as illustrated in Figure 1, where both covariate and concept shifts increase over time. Based on this observation, we hypothesize that an effective self-training approach can continuously adapt the model to account for these evolving shifts. Notably, prior work by Kumar et al. [43] establishes formal error bounds of self-training under the assumption of gradual distributional shifts. Our method builds on this foundation by extending their framework, and we provide a rigorous theoretical analysis in Appendix H.
@@ IV-A Adaptive Drift-Aware Pseudo-Labeling
Figure 3: Mean and standard deviation of predicted probabilities for benign and malware classes using XGBoost and MLP. Lower confidence for malware highlights that drift mainly affects this class.
@@ IV-B Data Augmentation with Label Consistency
[…]
Figure 4: Percentage of Drebin and APIGraph samples where augmentation flips a correct prediction (i.e., benign to malicious or vice versa).
@@
To address this, we propose two key modifications to ensure label consistency during augmentation. First, when replacing a masked feature, we restrict the replacement values only from samples of the same class. Although this setting was considered in [76], we argue that preserving label consistency is even more critical in our iterative training framework. Second, after the initial training phase, we leverage the current model to predict the label of the augmented sample. The augmented sample is included in the dataset only if the model’s prediction matches the original label to improve label consistency. However, it is important to note that feature consistency is not explicitly enforced during augmentation. This means the generated features may not always represent valid samples, potentially leading to inconsistencies.
@@ IV-C Confidence Calibration with Mixup
However, this approach is not applicable to models like Random Forests, which do not support fractional targets for classification tasks. To handle this, when mixing two samples with different labels, we assign the label corresponding to the sample with the higher interpolation coefficient. We provide evidence of how mixup improves confidence calibration in Section VII-D.
@@ VI Experimental Settings
TABLE II: Comparison of methods across all datasets. Results are the mean of five runs (five different random seeds). Rows labeled Oracle represent an ideal scenario where retraining uses ground-truth labels; these are provided for reference only and are not considered in best result highlighting. For each dataset, underlined F1-scores indicate the best baseline (excluding proposed and Oracle methods), and bold+underline values indicate the best overall result (including proposed methods).
Drebin APIGraph BODMAS EMBER PDF Malware Method Model F1 FPR FNR F1 FPR FNR F1 FPR FNR F1 FPR FNR F1 FPR FNR Offline ML RF 33.2 0.14 78.7 46.8 0.34 66.9 96.3 0.21 6.69 86.3 3.47 21.7 98.0 0.07 3.31 XGBoost 44.9 0.66 65.6 66.5 0.97 42.7 99.2 0.30 1.23 92.3 3.58 11.4 97.9 0.14 3.17 MLP 40.9 0.57 69.8 57.0 1.67 49.9 98.2 0.64 2.89 88.1 5.66 17.1 97.4 0.11 4.26 Oracle RF 53.4 0.12 61.3 78.9 0.24 31.8 96.6 0.19 6.22 87.1 3.38 20.4 98.8 0.08 1.85 XGBoost 68.3 0.52 41.2 85.5 0.67 17.6 99.6 0.21 0.61 93.1 3.46 10.1 99.2 0.09 1.04 MLP 74.6 0.53 33.1 89.0 0.45 14.4 98.8 0.50 1.77 90.0 5.16 14.2 97.6 0.09 3.87 Self-Training ARF 34.0 9.88 45.8 61.1 5.28 25.0 97.5 2.13 2.28 69.7 91.8 1.78 29.7 0.01 80.0 DE++ 34.7 7.84 47.1 42.3 0.33 65.6 90.8 15.1 1.83 76.9 24.9 22.8 60.8 19.3 0.28 Insomnia 49.8 0.70 60.6 68.1 2.05 31.8 97.2 0.56 4.81 87.1 6.06 18.5 88.2 0.09 15.4 MORSE 41.2 4.50 54.7 75.3 1.82 22.2 98.8 0.48 1.73 89.8 6.34 13.7 97.4 0.12 4.16 Proposed RF+ADAPT 49.8 0.21 64.3 67.2 0.34 46.0 96.1 0.22 7.11 86.2 3.56 21.7 99.1 0.09 1.26 XGBoost+ADAPT 57.1 0.24 56.6 73.2 2.26 22.2 99.6 0.24 0.61 93.5 4.45 8.70 98.6 0.10 2.03 MLP+ADAPT 50.5 0.77 59.1 76.8 0.81 29.4 98.6 0.41 2.26 89.1 5.01 16.0 97.2 0.09 4.73
@@ VI-B Self-Training Baselines
We compare our method against four self-training baselines: Adaptive Random Forest (ARF) [35], DroidEvolver++ (DE++) [41], Insomnia [9], and MORSE [76]. ARF and DE++ are online learning methods that address concept drift without requiring access to the original training dataset. In contrast, Insomnia and MORSE follow a semi-supervised paradigm, leveraging the original training data and pseudo-labeled samples for model updates. While our approach (ADAPT) also assumes access to the original training data, we additionally explore scenarios where this access is restricted. Further implementation details for all baseline methods are provided in Appendix A, and the restricted setting without original training data is discussed in Appendix D.
@@ VI-C Hyperparameter Tuning
As shown in previous work [24, 2], continuous learning settings require dedicated hyperparameter tuning. In our experiments, we jointly tune the hyperparameters of both the baseline models and our proposed approach using the validation set. ADAPT introduces five key hyperparameters: the benign base threshold ($\tau_{b}$), the malware base threshold ($\tau_{m}$), the adaptive threshold weight ($\lambda$), the masking probability for augmentation ($p_{a}$), and the mixup parameter ($\alpha$). Each baseline self-training method also has its hyperparameters; for example, ARF uses a threshold ($\tau$), DE++ tunes the app buffer size, malware-to-benign ratio, and model aging threshold, Insomnia adjusts the pseudo-labeling fraction, and MORSE tunes the threshold along with weak and strong augmentation probabilities. A complete list of hyperparameters and their ranges is provided in Appendix G. To ensure fair comparison despite varying hyperparameter counts, we follow [2] and use a fixed budget of 200 random hyperparameter searches [16] for each method on each dataset.